# HEIGHT: Heterogeneous Interaction Graph Transformer for Robot Navigation in Crowded and Constrained Environments

### v3

@@ IV-B3 Training
We train the entire network with Proximal Policy Optimization (PPO) [schulman2017proximal] in a simulator as shown in Fig. 2(a) (see Sec. V-A for details of the simulator). At each timestep $t$, the simulator provides all state information that constitutes $s^{t}$, which is fed to the HEIGHT network. The network outputs the estimated value of the state $V(s^{t})$ and the logarithmic probabilities of the robot’s action $\pi(a^{t}|s^{t})$, both of which are used to compute the PPO loss and then update the parameters in the network. In training, the robot action is sampled from the action distribution $\pi(a^{t}|s^{t})$. In testing, the robot takes the action with the highest probability $a^{t}$. The robot action $a^{t}$ is fed into the simulator to compute the next state $s^{t+1}$, and then the loop continues.
Without any supervised learning, our method is not limited by the performance of expert demonstrations [chen2017decentralized, chen2019crowd, liu2024sample]. However, to improve the low training data efficiency, an inherent problem of RL, HEIGHT can also be trained with a combination of imitation learning and RL.
@@ V-A Simulation benchmark
We conduct simulation experiments in a PyBullet simulator [coumans2019]. The robot, humans, and static obstacles are in a $\times12 m$ arena. In each episode, rectangular obstacles are initialized with random shapes and random poses. The width and the length of each static obstacle are sampled from $\mathcal{N}(1,0.6^{2})$ and are limited in $[0.1,5]$ meters. The initial positions of the humans and the robot are also randomized. The starting and goal positions of the robot are randomly sampled inside the arena. The distance between the robot’s starting and the goal positions is between 5 m and 6 m. Some humans are dynamic and some are static. The goals of dynamic humans are set on the opposite side of their initial positions to create circle-crossing scenarios. In training, the number of humans varies from 5 to 9 and the number of static rectangular obstacles varies from 8 to 12. Among all humans, 0-2 of them are static and the rest are dynamic. In testing, the number of humans and obstacles are shown in the first column of Table II. The heights of humans and obstacles are all above the height of the robot LiDAR to ensure detectability.
To simulate a continuous human flow similar to [liu2020decentralized, liu2023intention, xie2023drlvo], dynamic humans will move to new random goals immediately after they arrive at their goal positions or they get stuck in front of narrow passageways for more than 10 timesteps. All dynamic humans are controlled by ORCA [van2011reciprocal]. 80% of dynamic humans do not react to the robot and 20% of humans react to the robot. This mixed setting prevents our network from learning an extremely aggressive policy in which the robot forces all humans to yield while achieving a high reward. Meanwhile, the simulation maintains enough reactive humans to resemble the real crowd behaviors. We use holonomic kinematics for humans. The preferred speed of humans ranges from 0.4 m/s to 0.6 m/s to accommodate the speed of the robot. We assume that humans can achieve the desired velocities immediately, and they will keep moving with these velocities for the next $\Delta t$ seconds. The simulation and control frequency $\Delta t$ is 0.1 s. The maximum length of an episode is 491 timesteps or 49.1 s.
@@ V-B1 Baselines
DWA [fox1997dynamic] Circles for both Treats humans as obstacles, no interaction modeling
ORCA [van2011reciprocal] Circles for humans + Line segments for obstacles Optimization program with velocity obstacles to model all interactions
$\textrm{A}^{*}$+ CNN [perez2021robot] LiDAR point cloud for both A* + RL policy with no interaction modeling
DS-RNN [liu2020decentralized] Circles for both RL policy with only RH attention and GRU
DRL-VO [xie2023drlvo] Occupancy map for humans + Point cloud for both Pure pursuit algorithm + RL policy with no interaction modeling
HomoGAT [velickovic2018graph] Circles for humans + Point cloud for obstacles RL policy from st-graph with shared attention weights for all interactions
@@ V-B4 Evaluation
Training distribution 5-9 humans 8-12 obstacles DWA [fox1997dynamic] 0.16 0.68 0.63 0.05 0.15 28.45 11.52
ORCA [van2011reciprocal] 0.34 0.62 0.31 0.32 0.04 21.51 9.46
$\textrm{A}^{*}$+ CNN [perez2021robot] 0.64 0.29 0.28 0.01 0.07 25.72 12.30
DS-RNN [liu2020decentralized] 0.80 0.20 0.11 0.09 0.00 19.90 10.78
DRL-VO [xie2023drlvo] 0.59 0.41 0.34 0.07 0.00 21.45 10.36
HomoGAT [velickovic2018graph] 0.86 0.14 0.13 0.01 0.00 18.66 10.36
@@
Less crowded 0-4 humans 8-12 obstacles DWA [fox1997dynamic] 0.29 0.37 0.28 0.09 0.34 25.74 14.42
ORCA [van2011reciprocal] 0.42 0.54 0.11 0.42 0.04 19.77 9.69
$\textrm{A}^{*}$+ CNN [perez2021robot] 0.84 0.12 0.11 0.01 0.04 22.64 11.73
DS-RNN [liu2020decentralized] 0.72 0.28 0.07 0.21 0.00 17.21 9.56
DRL-VO [xie2023drlvo] 0.82 0.17 0.12 0.05 0.01 19.14 10.53
HomoGAT [velickovic2018graph] 0.95 0.05 0.03 0.02 0.00 16.64 10.17
@@
More crowded 10-14 humans 8-12 obstacles DWA [fox1997dynamic] 0.11 0.78 0.74 0.04 0.11 29.60 10.37
ORCA [van2011reciprocal] 0.26 0.71 0.41 0.30 0.03 23.33 9.21
$\textrm{A}^{*}$+ CNN [perez2021robot] 0.47 0.42 0.39 0.03 0.11 27.33 12.47
DS-RNN [liu2020decentralized] 0.76 0.22 0.16 0.06 0.01 23.08 11.67
DRL-VO [xie2023drlvo] 0.50 0.49 0.41 0.09 0.01 22.72 10.26
HomoGAT [velickovic2018graph] 0.72 0.28 0.23 0.05 0.00 20.46 10.57
@@
Less constrained 5-9 humans 3-7 obstacles DWA [fox1997dynamic] 0.17 0.57 0.55 0.02 0.26 28.50 14.21
ORCA [van2011reciprocal] 0.42 0.54 0.30 0.24 0.03 21.89 10.59
$\textrm{A}^{*}$+ CNN [perez2021robot] 0.66 0.29 0.28 0.01 0.05 23.63 11.98
DS-RNN [liu2020decentralized] 0.66 0.34 0.20 0.14 0.00 18.01 9.76
DRL-VO [xie2023drlvo] 0.70 0.30 0.26 0.04 0.00 21.21 11.15
HomoGAT [velickovic2018graph] 0.88 0.12 0.10 0.02 0.00 17.66 10.37
@@
More constrained 5-9 humans 13-17 obstacles DWA [fox1997dynamic] 0.14 0.66 0.49 0.17 0.20 30.47 11.35
ORCA [van2011reciprocal] 0.30 0.58 0.24 0.34 0.11 22.40 9.04
$\textrm{A}^{*}$+ CNN [perez2021robot] 0.48 0.29 0.23 0.06 0.23 27.28 13.11
DS-RNN [liu2020decentralized] 0.71 0.23 0.09 0.14 0.06 23.45 11.58
DRL-VO [xie2023drlvo] 0.55 0.40 0.23 0.07 0.05 21.87 10.22
HomoGAT [velickovic2018graph] 0.80 0.20 0.11 0.09 0.00 19.20 10.51
@@ V-C4 Failure cases
By visualizing the testing episodes across all environments, we find that HEIGHT typically collides when (1) a human arrives at its goal and suddenly changes directions to a new goal (Fig. 11(a) in Appendix) or (2) the robot surroundings are extremely crowded and constrained, and almost all free paths are blocked by humans (Fig. 11(b)). In these cases, due to its speed limit, the robot sometimes cannot switch to an alternative path in time to prevent collisions. To remedy the difficulty of RL in long-term decision making, in future work, our method can be combined with motion primitives and path planning algorithms that consider the stochasticity of pedestrian motions [hu2025composablenav].
@@ VI-A Setup
In the Hallway and Lounge environments, we conduct controlled experiments where the trajectories of all pedestrians are fixed to ensure a fair comparison. We compare our method with the following two baselines: (1) ROS navigation stack is a geometric planner which uses the dynamic window approach (DWA) [fox1997dynamic] as the local planner and $\text{A}^{*}$ as the global planner; (2) DRL-VO is a state-of-the-art RL method [xie2023drlvo]. For each method, we run 30 trials in total. Among all trials, the start goal positions of the robot are fixed. We measure the success, collision, timeout rates, and navigation time of successful episodes as testing metrics.
@@ VII-A Sim2real through Real2sim
Limitation and future work: Although we have minimized the sim2real gaps through real2sim, certain gaps still exist. Before sim2real transfer, the policy needs to be finetined in simulations that mimic the target environments (Fig.9). To further align the simulation and the real-world, we plan to explore the following directions for future work: (1) developing a more natural pedestrian model to replace the ORCA humans in the simulator, (2) revising our pipeline to enable failure detections and self-supervised RL fine-tuning in the real-world [Zhu2020The, chang2023data], (3) using a small amount of real-world data to automatically optimize the parameters of our simulator to match the real-world environment [du2021auto, lim2022real2sim2real].
@@ VII-B Scene representation
Limitation and future work: A side effect of our scene abstraction is the loss of detailed information such as gaits of humans and 3D shapes of obstacles. Another side effect is the cascading errors between the perception modules and robot policy, such as inaccurate human detections or robot localization. A future direction is the joint optimization of the whole robotic stack from perception to control [hu2023_uniad, wang2024mosaic].
@@ VII-C Structured neural network
Limitation and future work: The structured neural network utilizes the inductive bias behind human behaviors to solve specific problems. However, without re-training, the network cannot generalize to other tasks such as autonomous driving. In other fields such as computer vision and natural language processing, foundation models that are trained on large-scale diverse data have shown promise in generalization. Inspired by the standardized benchmarks of autonomous driving and robot manipulation [caesar2021nuplan, openxembodiment2023], we aim to contribute to the benchmarking efforts of human-centered tasks.
@@ VII-D Training method
Limitation and future work: Our method is subjective to the inherent limitations of RL, such as the need for simulators, low training data efficiency, and difficulties in long-horizon tasks. In addition, preventing performance degradation of the policy when inevitable distribution shifts happen, especially in real-world, is challenging. In future work, combining the RL policy with other planners, imitation learning, or foundation models may alleviate these problems [raj2024rethinking].
@@ VII-E Miscellaneous
Social awareness: The attention-based interaction framework presents a promising foundation for some social behaviors (e.x. treating humans and obstacles differently). However, HEIGHT focuses on optimizing task success and efficiency, and thus lacks explicit mechanisms and evaluation metrics for social awareness [francis2025principles]. Future work could integrate social norm as constraints or auxiliary objectives in RL training, as well as learning from human demonstrations or preference-based feedback. User studies can evaluate social awareness.
@@ Acknowledgements
  • Dynamic window approach (DWA) [fox1997dynamic] searches for the optimal velocity that brings the robot to the goal with the maximum clearance from any obstacle. DWA is a model-based method that only considers the current state. DWA represents both humans and static obstacles as groups of small circles. We re-mapped the output of DWA to match the turtlebot kinematics and our action space.
  • Optimal Reciprocal Collision Avoidance (ORCA) [van2011reciprocal] optimizes for the optimal velocity that brings the robot to the goal by assuming that agents avoid each other under the reciprocal rule. ORCA is a model-based method that only considers the current state. In addition, ORCA represents both humans as circles and static obstacles as line segments. We re-mapped the output of ORCA to match the unicycle kinematics of our robot.
  • $\textrm{A}^{*}$+ CNN [perez2021robot] is a hybrid method with sample-based planning and RL. With a map of the environment, $\textrm{A}^{*}$ is the global planner and generates 6 waypoints in the beginning of an episode. The inputs to the RL policy are a 2D LiDAR point cloud, the robot state, the waypoints, and the robot goal. No human detections are used and human features are included in the point cloud. The point cloud is passed through a CNN, whose output is concatenated with the embedding of other inputs. In addition to the reward function in Eq. 1, the robot receives a heading direction award if it moves towards the goal while being collision-free.
  • Decentrialized structural RNN (DS-RNN) [liu2020decentralized] is an RL-based method that represents static obstacles as groups of small circles. In network architecture, DS-RNN only contains the RH attention and the GRU.
  • DRL-VO [xie2023drlvo] is a hybrid method with a pure pursuit algorithm as the global planner and an RL policy as the local planner. The inputs to the planners are the positions and velocities of humans encoded in an occupancy map, a 2D LiDAR point cloud, and the robot state. The RL network is a CNN without explicit interaction modeling.
  • Homogeneous graph attention network (HomoGAT) [velickovic2018graph] is RL-based and splits human and obstacle inputs. However, HomoGAT does not differentiate between 3 types of nodes and 3 types of edges in policy network. Instead, HomoGAT uses a single self-attention network to weigh humans, the robot, and the obstacle point cloud and feed the weighted sum of all embeddings into a GRU. HomoGAT is similar to Chen et al. [chen2019relational] and Liu et al. [liu2024sample] but the input, output, and training algorithm are kept the same as our method for a fair comparison.
@@
DWA [fox1997dynamic] N/A 1.5
ORCA [van2011reciprocal] N/A 0.3
$\textrm{A}^{*}$+ CNN [perez2021robot] 47.2 2.0
DS-RNN [liu2020decentralized] 49.0 1.0
DRL-VO [xie2023drlvo] 111.9 2.3
HomoGAT [velickovic2018graph] 47.7 2.2
@@
  • In Hallway and Lounge environments, we use a TurtleBot 2i equipped with an RPLIDAR-A3 laser scanner, an Intel RealSense tracking camera T265, a Nvidia Jetson Xaiver board, and a remote host computer. We first use the ROS gmapping package to create a map of the environment. Then, we process the map to combine small obstacles and eliminate noises. To detect human positions from LiDAR point clouds, we use an off-the-shelf people detector, DR-SPAAM [Jia2020DRSPAAM]. Since DR-SPAAM tends to treat obstacles as false positive humans, we mitigate this issue by removing the point clouds that belong to pre-mapped static obstacles from the LiDAR readings, then feed the remaining point clouds that only contain humans to DR-SPAAM. We use the T265 camera to obtain the pose of the robot $(p_{x},p_{y},\theta)$ and the robot wheel encoder to obtain its velocity $(u_{x},u_{y})$. The host computer runs the robot policies and perception modules, while the TurtleBot only receives and sends ROS messages. The host computer and TurtleBot communicate through ROS.
  • In Atrium and Outdoor environment, we use a Clearpath Jackal robot equipped with an ZED 2i camera and an Ouster LiDAR. The system is built using ROS. The software stack is similar to the TurtleBot except the following differences: We use ZED 2i to detect the positions of humans and the pose of the robot. The pre-built map and robot localization are obtained from ENML algorithm [enml]. All computations are executed entirely onboard, utilizing an Intel i7-9700TE CPU and an NVIDIA RTX A2000 GPU.

### v4

@@ IV-B3 Training
We train the entire network with Proximal Policy Optimization (PPO) [55] in a simulator as shown in Fig. 2(a) (see Sec. V-A for details of the simulator). At each timestep $t$, the simulator provides all state information that constitutes $s^{t}$, which is fed to the HEIGHT network. The network outputs the estimated value of the state $V(s^{t})$ and the logarithmic probabilities of the robot’s action $\pi(a^{t}|s^{t})$, both of which are used to compute the PPO loss and then update the parameters in the network. In training, the robot action is sampled from the action distribution $\pi(a^{t}|s^{t})$. In testing, the robot takes the action with the highest probability $a^{t}$. The robot action $a^{t}$ is fed into the simulator to compute the next state $s^{t+1}$, and then the loop continues.
Without any supervised learning, our method is not limited by the performance of expert demonstrations [9, 7, 25]. However, to improve the low training data efficiency, an inherent problem of RL, HEIGHT can also be trained with a combination of imitation learning and RL.
@@ V-A Simulation benchmark
We conduct simulation experiments in a PyBullet simulator [12]. The robot, humans, and static obstacles are in a $\times12 m$ arena. In each episode, rectangular obstacles are initialized with random shapes and random poses. The width and the length of each static obstacle are sampled from $\mathcal{N}(1,0.6^{2})$ and are limited in $[0.1,5]$ meters. The initial positions of the humans and the robot are also randomized. The starting and goal positions of the robot are randomly sampled inside the arena. The distance between the robot’s starting and the goal positions is between 5 m and 6 m. Some humans are dynamic and some are static. The goals of dynamic humans are set on the opposite side of their initial positions to create circle-crossing scenarios. In training, the number of humans varies from 5 to 9 and the number of static rectangular obstacles varies from 8 to 12. Among all humans, 0-2 of them are static and the rest are dynamic. In testing, the number of humans and obstacles are shown in the first column of Table II. The heights of humans and obstacles are all above the height of the robot LiDAR to ensure detectability.
To simulate a continuous human flow similar to [36, 35, 67], dynamic humans will move to new random goals immediately after they arrive at their goal positions or they get stuck in front of narrow passageways for more than 10 timesteps. All dynamic humans are controlled by ORCA [59]. 80% of dynamic humans do not react to the robot and 20% of humans react to the robot. This mixed setting prevents our network from learning an extremely aggressive policy in which the robot forces all humans to yield while achieving a high reward. Meanwhile, the simulation maintains enough reactive humans to resemble the real crowd behaviors. We use holonomic kinematics for humans. The preferred speed of humans ranges from 0.4 m/s to 0.6 m/s to accommodate the speed of the robot. We assume that humans can achieve the desired velocities immediately, and they will keep moving with these velocities for the next $\Delta t$ seconds. The simulation and control frequency $\Delta t$ is 0.1 s. The maximum length of an episode is 491 timesteps or 49.1 s.
@@ V-B1 Baselines
DWA [17] Circles for both Treats humans as obstacles, no interaction modeling
ORCA [59] Circles for humans + Line segments for obstacles Optimization program with velocity obstacles to model all interactions
$\textrm{A}^{*}$+ CNN [48] LiDAR point cloud for both A* + RL policy with no interaction modeling
DS-RNN [36] Circles for both RL policy with only RH attention and GRU
DRL-VO [67] Occupancy map for humans + Point cloud for both Pure pursuit algorithm + RL policy with no interaction modeling
HomoGAT [62] Circles for humans + Point cloud for obstacles RL policy from st-graph with shared attention weights for all interactions
@@ V-B4 Evaluation
Training distribution 5-9 humans 8-12 obstacles DWA [17] 0.16 0.68 0.63 0.05 0.15 28.45 11.52
ORCA [59] 0.34 0.62 0.31 0.32 0.04 21.51 9.46
$\textrm{A}^{*}$+ CNN [48] 0.64 0.29 0.28 0.01 0.07 25.72 12.30
DS-RNN [36] 0.80 0.20 0.11 0.09 0.00 19.90 10.78
DRL-VO [67] 0.59 0.41 0.34 0.07 0.00 21.45 10.36
HomoGAT [62] 0.86 0.14 0.13 0.01 0.00 18.66 10.36
@@
Less crowded 0-4 humans 8-12 obstacles DWA [17] 0.29 0.37 0.28 0.09 0.34 25.74 14.42
ORCA [59] 0.42 0.54 0.11 0.42 0.04 19.77 9.69
$\textrm{A}^{*}$+ CNN [48] 0.84 0.12 0.11 0.01 0.04 22.64 11.73
DS-RNN [36] 0.72 0.28 0.07 0.21 0.00 17.21 9.56
DRL-VO [67] 0.82 0.17 0.12 0.05 0.01 19.14 10.53
HomoGAT [62] 0.95 0.05 0.03 0.02 0.00 16.64 10.17
@@
More crowded 10-14 humans 8-12 obstacles DWA [17] 0.11 0.78 0.74 0.04 0.11 29.60 10.37
ORCA [59] 0.26 0.71 0.41 0.30 0.03 23.33 9.21
$\textrm{A}^{*}$+ CNN [48] 0.47 0.42 0.39 0.03 0.11 27.33 12.47
DS-RNN [36] 0.76 0.22 0.16 0.06 0.01 23.08 11.67
DRL-VO [67] 0.50 0.49 0.41 0.09 0.01 22.72 10.26
HomoGAT [62] 0.72 0.28 0.23 0.05 0.00 20.46 10.57
@@
Less constrained 5-9 humans 3-7 obstacles DWA [17] 0.17 0.57 0.55 0.02 0.26 28.50 14.21
ORCA [59] 0.42 0.54 0.30 0.24 0.03 21.89 10.59
$\textrm{A}^{*}$+ CNN [48] 0.66 0.29 0.28 0.01 0.05 23.63 11.98
DS-RNN [36] 0.66 0.34 0.20 0.14 0.00 18.01 9.76
DRL-VO [67] 0.70 0.30 0.26 0.04 0.00 21.21 11.15
HomoGAT [62] 0.88 0.12 0.10 0.02 0.00 17.66 10.37
@@
More constrained 5-9 humans 13-17 obstacles DWA [17] 0.14 0.66 0.49 0.17 0.20 30.47 11.35
ORCA [59] 0.30 0.58 0.24 0.34 0.11 22.40 9.04
$\textrm{A}^{*}$+ CNN [48] 0.48 0.29 0.23 0.06 0.23 27.28 13.11
DS-RNN [36] 0.71 0.23 0.09 0.14 0.06 23.45 11.58
DRL-VO [67] 0.55 0.40 0.23 0.07 0.05 21.87 10.22
HomoGAT [62] 0.80 0.20 0.11 0.09 0.00 19.20 10.51
@@ V-C4 Failure cases
By visualizing the testing episodes across all environments, we find that HEIGHT typically collides when (1) a human arrives at its goal and suddenly changes directions to a new goal (Fig. 11(a) in Appendix) or (2) the robot surroundings are extremely crowded and constrained, and almost all free paths are blocked by humans (Fig. 11(b)). In these cases, due to its speed limit, the robot sometimes cannot switch to an alternative path in time to prevent collisions. To remedy the difficulty of RL in long-term decision making, in future work, our method can be combined with motion primitives and path planning algorithms that consider the stochasticity of pedestrian motions [24].
@@ VI-A Setup
In the Hallway and Lounge environments, we conduct controlled experiments where the trajectories of all pedestrians are fixed to ensure a fair comparison. We compare our method with the following two baselines: (1) ROS navigation stack is a geometric planner which uses the dynamic window approach (DWA) [17] as the local planner and $\text{A}^{*}$ as the global planner; (2) DRL-VO is a state-of-the-art RL method [67]. For each method, we run 30 trials in total. Among all trials, the start goal positions of the robot are fixed. We measure the success, collision, timeout rates, and navigation time of successful episodes as testing metrics.
@@ VII-A Sim2real through Real2sim
Limitation and future work: Although we have minimized the sim2real gaps through real2sim, certain gaps still exist. Before sim2real transfer, the policy needs to be finetined in simulations that mimic the target environments (Fig.9). To further align the simulation and the real-world, we plan to explore the following directions for future work: (1) developing a more natural pedestrian model to replace the ORCA humans in the simulator, (2) revising our pipeline to enable failure detections and self-supervised RL fine-tuning in the real-world [72, 4], (3) using a small amount of real-world data to automatically optimize the parameters of our simulator to match the real-world environment [14, 33].
@@ VII-B Scene representation
Limitation and future work: A side effect of our scene abstraction is the loss of detailed information such as gaits of humans and 3D shapes of obstacles. Another side effect is the cascading errors between the perception modules and robot policy, such as inaccurate human detections or robot localization. A future direction is the joint optimization of the whole robotic stack from perception to control [23, 64].
@@ VII-C Structured neural network
Limitation and future work: The structured neural network utilizes the inductive bias behind human behaviors to solve specific problems. However, without re-training, the network cannot generalize to other tasks such as autonomous driving. In other fields such as computer vision and natural language processing, foundation models that are trained on large-scale diverse data have shown promise in generalization. Inspired by the standardized benchmarks of autonomous driving and robot manipulation [2, 11], we aim to contribute to the benchmarking efforts of human-centered tasks.
@@ VII-D Training method
Limitation and future work: Our method is subjective to the inherent limitations of RL, such as the need for simulators, low training data efficiency, and difficulties in long-horizon tasks. In addition, preventing performance degradation of the policy when inevitable distribution shifts happen, especially in real-world, is challenging. In future work, combining the RL policy with other planners, imitation learning, or foundation models may alleviate these problems [51].
@@ VII-E Miscellaneous
Social awareness: The attention-based interaction framework presents a promising foundation for some social behaviors (e.x. treating humans and obstacles differently). However, HEIGHT focuses on optimizing task success and efficiency, and thus lacks explicit mechanisms and evaluation metrics for social awareness [18]. Future work could integrate social norm as constraints or auxiliary objectives in RL training, as well as learning from human demonstrations or preference-based feedback. User studies can evaluate social awareness.
@@ Acknowledgements
  • Dynamic window approach (DWA) [17] searches for the optimal velocity that brings the robot to the goal with the maximum clearance from any obstacle. DWA is a model-based method that only considers the current state. DWA represents both humans and static obstacles as groups of small circles. We re-mapped the output of DWA to match the turtlebot kinematics and our action space.
  • Optimal Reciprocal Collision Avoidance (ORCA) [59] optimizes for the optimal velocity that brings the robot to the goal by assuming that agents avoid each other under the reciprocal rule. ORCA is a model-based method that only considers the current state. In addition, ORCA represents both humans as circles and static obstacles as line segments. We re-mapped the output of ORCA to match the unicycle kinematics of our robot.
  • $\textrm{A}^{*}$+ CNN [48] is a hybrid method with sample-based planning and RL. With a map of the environment, $\textrm{A}^{*}$ is the global planner and generates 6 waypoints in the beginning of an episode. The inputs to the RL policy are a 2D LiDAR point cloud, the robot state, the waypoints, and the robot goal. No human detections are used and human features are included in the point cloud. The point cloud is passed through a CNN, whose output is concatenated with the embedding of other inputs. In addition to the reward function in Eq. 1, the robot receives a heading direction award if it moves towards the goal while being collision-free.
  • Decentrialized structural RNN (DS-RNN) [36] is an RL-based method that represents static obstacles as groups of small circles. In network architecture, DS-RNN only contains the RH attention and the GRU.
  • DRL-VO [67] is a hybrid method with a pure pursuit algorithm as the global planner and an RL policy as the local planner. The inputs to the planners are the positions and velocities of humans encoded in an occupancy map, a 2D LiDAR point cloud, and the robot state. The RL network is a CNN without explicit interaction modeling.
  • Homogeneous graph attention network (HomoGAT) [62] is RL-based and splits human and obstacle inputs. However, HomoGAT does not differentiate between 3 types of nodes and 3 types of edges in policy network. Instead, HomoGAT uses a single self-attention network to weigh humans, the robot, and the obstacle point cloud and feed the weighted sum of all embeddings into a GRU. HomoGAT is similar to Chen et al. [6] and Liu et al. [25] but the input, output, and training algorithm are kept the same as our method for a fair comparison.
@@
DWA [17] N/A 1.5
ORCA [59] N/A 0.3
$\textrm{A}^{*}$+ CNN [48] 47.2 2.0
DS-RNN [36] 49.0 1.0
DRL-VO [67] 111.9 2.3
HomoGAT [62] 47.7 2.2
@@
  • In Hallway and Lounge environments, we use a TurtleBot 2i equipped with an RPLIDAR-A3 laser scanner, an Intel RealSense tracking camera T265, a Nvidia Jetson Xaiver board, and a remote host computer. We first use the ROS gmapping package to create a map of the environment. Then, we process the map to combine small obstacles and eliminate noises. To detect human positions from LiDAR point clouds, we use an off-the-shelf people detector, DR-SPAAM [29]. Since DR-SPAAM tends to treat obstacles as false positive humans, we mitigate this issue by removing the point clouds that belong to pre-mapped static obstacles from the LiDAR readings, then feed the remaining point clouds that only contain humans to DR-SPAAM. We use the T265 camera to obtain the pose of the robot $(p_{x},p_{y},\theta)$ and the robot wheel encoder to obtain its velocity $(u_{x},u_{y})$. The host computer runs the robot policies and perception modules, while the TurtleBot only receives and sends ROS messages. The host computer and TurtleBot communicate through ROS.
  • In Atrium and Outdoor environment, we use a Clearpath Jackal robot equipped with an ZED 2i camera and an Ouster LiDAR. The system is built using ROS. The software stack is similar to the TurtleBot except the following differences: We use ZED 2i to detect the positions of humans and the pose of the robot. The pre-built map and robot localization are obtained from ENML algorithm [1]. All computations are executed entirely onboard, utilizing an Intel i7-9700TE CPU and an NVIDIA RTX A2000 GPU.